Thomas is single and a self-employed architect. During 2019, Thomas’s income from his business is $150,000. He also pays $2,180 for a medical insurance policy.

What is his 2019 self-employment tax deduction? Round intermediate calculations and final answers to the nearest whole dollar. Net self-employment income is total income multiplied by 92.35%. Of this amount, 15.3% is the self-employment tax. The deduction is one-half of the amount for the self-employment tax. Also, remember that OASDI is taxed on a maximum of $132,900 and MHI of 2.9% is taxed on the total net self-employment income.

What is his self-employment tax in 2019?

What is his deduction for adjusted gross income in 2019?

Assume that Thomas's income from his business is $234,000. What is his 2019 self-employment tax deduction? Round intermediate calculations and final answers to the nearest whole dollar. HI is taxed at 0.9% on the excess of net self-employment income over $200,000.

What is his self-employment tax in 2019?

What is his deduction for adjusted gross income in 2019?
Answer #1

Part A

Total self-employment tax

$20497

Deduction for adjusted gross income

$10249

Net Self-employment income = 150000*92.35% = 138525

Self-employment income is over the ceiling of $132900 in 2019,the OASDI tax of 12.4% (6.2% x 2) on $132900 and MHI tax of 2.9% (1.45% x 2) on $138525.

Net Self-employment income

138525

OASDI (132900*12.4%)

16480

MHI (138525*2.9%)

4017

Total self-employment tax

$20497

Deduction for adjusted gross income (20497*50%)

$10249

Part B

Total self-employment tax

$22892

Deduction for adjusted gross income

$11446

Net Self-employment income = 234000*92.35% = 216099

Self-employment income is over the ceiling of $132900 in 2019,the OASDI tax of 12.4% (6.2% x 2) on $132900 and MHI tax of 2.9% (1.45% x 2) on $216099. As Net Self-employment income exceeds $200000, additional tax of 0.9% is applied on $16099 (216099-200000).

Net Self-employment income

216099

OASDI (132900*12.4%)

16480

MHI (216099*2.9%)

6267

HI (16099*0.9%)

145

Total self-employment tax

$22892

Deduction for adjusted gross income (22892*50%)

$11446